73 C CO DOO drbd d O wnu codspoq 6 9 uped dnbd awoy busds asnod SnS Neas o aay ada WYLEO 1siawsue nol posos o O o au pue uo paedu m q u aouaagp e s aa prpu3 4 pa p as O Moja asue paio a asooND adoudde wg aprs (papaau se upuesnoa yaeau ag o puno) punog addn au st punog m gnou sugu pagedu, se pandauoo ae seouaap a a uop ou pue uop pagedu mou f u aouaagp e s a s uappo 6sn (k yenous a ou g azs adues ag asneaq uauada u p u eape poobe p ae saug uogpeas asneoaq uada buubsap u rape poob buubiap u eape poob e s4 O s usn u ancoo e jeSuuea ue o ouoasveoq apada 41 VO ad Euusap u eap poob e sp wM papas uopuei sem yg sa6o6 aomo paedugn pey aprs o n(e) sagno ou panqasp geuou Ae adde ae soouagp ou ep apeop eep ap o yodsog pue yopd Augaegod peou yoN () pue (e) sued aaduog papucoau sEM noy ad sagu g9 po poods e wog dos e o apyan ou Buq o paanbau (spucoas u) ouug au saobeuo ouu o adues uopues e o uo uo jooe o spage a aus o sobob oads yo sed e Buesn opp ospe pue uos paeduun paansesu seM foouu ue u oug uogoeas juapns eoups o pus inae wnog go aod e oupoe o na) 0ung opun Bup o spoga oeng 01 da uogsano
5 (Round to the nearest thousandth as needed.) State the appropriate conclusion. Choose the correct answer below. There is insufficient evidence to conclude there is a difference in braking time with impaired vision and normal vision. There is sufficient evidence to conclude there is a difference in braking time with impaired vision and normal vision. Click to select your answer(s). Type here to search SV Myh Sonic Master 91 O/R 61 OLI X % 6
hol, a police officer brought a DUI simulator to a local high school. Student reaction time in an emerg on vision. Fora random sample of nine teenagers, the time (in seconds) required to bring the vehicl P allvdictributodwith the diffaren Data Table control action e samp 6 7 8 9 4.58 4.56 4.31 4.80 4.59 5.00 4.79 Subject 4 5 raking Normal, Xi 4.49 4.24 Impaired, Y 5.86 5.67 5.45 5.29 5.83 5.49 5.23 5.63 5.63 Print Done in braking time with impaired vision and normal vision. braking time with impaired vision and normal vision.
Answer #1

a) This is a good idea in designing the experiment because it controls for any “learning” that may occur in using the simulators.

b) 95% confidence interval :

First we need to find the differences (d) = impaired - Normal

Then we have to find the mean of differences \bar{d} and standard deviation of differences Sd using the excel function =AVERAGE() and STDEV.S() respectively.

A B C D Impaired Normal d 1 2 5.86 4.49 1.37 3 5.67 4.24 1.43 4 5.45 .58 -AVERAGE (C2 :C10) 0.9689 -STDEV.S(C2:C10) 5 5.29 4.

\bar{d} = 0.9689 and Sd = 0.3643

Lower bound =  phpVCNdxt.png - E

Upper bound =  phpeRoiNo.png + E

E = t* S Vn ; t is critical value follows t distribution with degrees of freedom (df) = n-1

We are given confidence level = 0.95

We can find critical value using excel function =TINV( α , d.f )

α = 1 - confidence level = 1 - 0.95 = 0.05

degrees of freedom (d.f) = n - 1 = 9-1 = 8

t =TINV(0.05,8)

t = 2.306

Therefore,

E = 2.3060.3643

E = 0.28

Lower bound =  phpyeppNh.png - E = 0.9689 - 0.28 = 0.689

Upper bound =  phpyXRfoO.png + E = 0.9689 + 0.28 = 1.249

95% Confidence interval is ( 0.689,1.249 )

Conclusion : As confidence interval does not contain 0 ;

There is sufficient evidence to conclude that there is difference in braking time with impaired vision and normal vision .